# Clotho: Measuring Task-Specific Pre-Generation Test Adequacy for LLM Inputs

### v2

@@ 5.2 Test Prioritisation Effectiveness (RQ2)
5.3 RQ3: Complementarity Analysis
RQ3 compares Clotho to post-generation approaches. Table 6 begins by comparing the results of correlation analysis, i.e., how well the adequacy scores correlate with pass rates, across Clotho and post-generation metrics: the highest are typeset in bold, and the second highest underlined. Post-generation variance based metrics generally perform well, with LOHSvar achieving the strongest correlations overall (6/8 for Llama, 5/8 for Mistral). In contrast, token-based confidence measures generally show only moderate correlations across all configurations. Overall, Clotho demonstrates strong performance, despite lacking access to the actual model outputs. Notably, Clotho achieves the best performance on specific tasks, particularly for SYN-BUG, where post-generation approaches typically underperform. We also note that the performance distribution varies significantly across model architectures. For Gemma, Clotho shows competitive performance compared to variance-based measures, suggesting particular compatibility with certain model families.
@@ 5.3 RQ3: Complementarity Analysis
Figure 7: 3-way Venn diagrams drawn from samples with high adequacy scores (top 25%) each from Clotho, LOHSvar, and TOKent among high-failure-rate inputs ($p_{f}\geq 0.8$) in Llama.
@@
Table 7: Spearman’s rank correlation between scores obtained from OLMs with reference set size 500 and actual pass rates of proprietary LLMs.
Gemma LLaMA Mistral
@@
5.4 RQ4: Transferability to Proprietary LLMs
The final research question focuses on the more practical applicability of Clotho: by design, Clotho requires internal representations of the target LLMs that are only available from Open-weight Language Models (OLMs). However, some application may have to rely on stronger, proprietary closed-weight models. Table 7 reports score correlation analysis results, in which the adequacy scores are computed by Clotho using open-weight models while the pass rates are obtained from 2,000 queries made to each of Claude, GPT, and Gemini. The results show that the adequacy scores predicted using OLMs can still produce moderate correlations with the actual pass rates of proprietary models: in some cases, even outperform the correlation against pass rates from the OLM itself (contained in the SELF column). For example, Gemma scores achieve higher correlation when applied to Gemini for MODEL-EX than when applied to itself; similarly, Mistral scores perform better when applied to GPT-4o mini then itself for POS-TAG task. This suggests that Clotho can capture certain generalisable aspects of input difficulty that transcend specific model architectures.
Figure 8: Average percentage of detected failures when scores from Open-weight Language Models (OLMs) are applied to proprietary LLMs.
We subsequently evaluate the effectiveness of these transferred adequacy scores for prioritizing failure-inducing inputs for the proprietary models. Figure 8 shows the cumulative failure rates for the transferred input orderings for GH-TYPO, JSON-FIX, SPELL-CHK, SYN-BUG, and ODD-ADD for a subset of transfer pairs (best transfer pair for each task): the solid lines denote the average from 10 runs of Clotho, and the coloured bands denote the standard deviation. We also report the Average Percentage of Fault Detected (APFD) Elbaum et al. (2000), which is essentially the area under curve. Given that average APFD would converge to 0.5 as the number of runs increases, the transferred ordering can achieve significantly effective test prioritisation. This is because proprietary models tend to have lower failure rate. As such, prioritising failure-inducing inputs early on can have bigger impact in proprietary models than in OLMs. Table 8 contains the failure@$N$ from transferred Clotho scores at NN=500, as well as random orderings, per task: the large gains over random ordering aligns with high APFDs we observe in Figure 8. Aggregated across tasks and models, Clotho finds on average 23.75 additional failures at NN=100 and 61.15 at NN=500 over random selection, corresponding to 126.8% and 64.5% improvements, respectively.
Table 8: Number of detected failing inputs at NN labelling budgets (failure@NN) when scores from OLMs are applied to proprietary model (NN=500).
Claude GPT Gemini
Gemma Llama Mistral Gemma Llama Mistral Gemma Llama Mistral
@@ 6 Discussion
6.1 Distinguishing Different Failure Modes
Our evaluation focuses on Clotho’s ability to identify inputs that trigger failures, aligning with the core aim of robustness testing: prioritizing failure-inducing cases. Since failures are relatively common in LLMs (as confirmed in our studied tasks), testers may also want to distinguish among types of failures. For example, in syntactic bug detection, it is often more valuable to gather one instance from several bug categories (e.g., missing brackets, incorrect keywords) than to collect many instances of the same type.
Since Clotho partitions the input space with a Gaussian Mixture Model (GMM), its components may reveal distinct modes of failure. Targeting low-likelihood inputs from different components could increase the diversity of failures found. Beyond this, building models of failing inputs, potentially one per failure category, alongside the passing-input model offers another path to sharpen failure discovery and characterisation.
@@ 6.1 Distinguishing Different Failure Modes
6.2 Tasks Beyond Binary Correctness
Tasks in our study, like many existing LLM benchmarks, adopt binary correctness criteria. Yet real-world applications often demand richer evaluation, with graded or multi-dimensional measures such as politeness, factual accuracy, or consistency in a counselling chatbot. We account for non-determinism through majority-pass criteria over repeated runs, but the idea of a “passing” input can naturally extend to thresholding continuous scores or aggregating grades across dimensions, as seen in recent evaluation frameworks PromptFoo (2024); AI (2024a, b); LangChain (2024). We hypothesise that Clotho can adapt to such settings by building reference sets around properties of interest and by defining adequacy relative to those properties. Exploring how thresholding and multi-criteria aggregation shape Clotho ’s performance will be essential to broadening its scope.
@@ 6.2 Tasks Beyond Binary Correctness
7 Threats to Validity
@@
Internal validity could be threatened because LLMs are not fully interpretable, meaning that Clotho might capture latent patterns that correlate with but are not identical to task-specific difficulty; using open-weight models may improve transparency and reproducibility but does not resolve this gap. External validity could also be at risk since Clotho learns task- and model-specific input distributions, which might limit generalisation, especially if other architectures such as reasoning LLMs (e.g., DeepSeek R1 DeepSeek-AI et al. (2025)) produce different latent spaces; we attempt to reduce this risk by evaluating multiple models and tasks, though broader studies will be useful. Construct validity could likewise be threatened if our chosen metrics do not fully reflect adequacy, and while we use established measures (ROC-AUC, APFD) and a scrutinised, open-source GMM implementation Pedregosa et al. (2011), these proxies may not perfectly represent the underlying concept.
8 Related Work
@@
Post-Generation Uncertainty: LLM outputs have been widely studied to measure uncertainty and, consequently, detect hallucinations. Chen et al. propose EigenScore, which leverages the eigenvalue decomposition of activation covariance matrices to detect hallucinations  Chen et al. (2024). The MIND approach of Su et al. goes further and toward online use, showing that unsupervised detectors over streaming internal states can flag hallucinations in real time during decoding Su et al. (2024). Factoscope from He et al. aggregates multi-layer inner states with a Siamese network architecture to separate factual from non-factual outputs across models, reinforcing that hidden-state geometry correlates with veracity He et al. (2024). Clotho, on the other hand, uses only the internal hidden space of inputs.
@@
Input-side Internal Activation: Recent work show that an LLM’s input-side internal activations already contain strong signals about errors in the output. Azaria and Mitchell train a lightweight classifier on hidden states to distinguish true statements from the false, outperforming probability-only baselines and demonstrating that pre-generation representations encode truthfulness cues Azaria and Mitchell (2023). Snyder et al. probe tokens and activations before the hallucination begins and train binary detectors that can forecast hallucinations in factual QA Snyder et al. (2024). Extending this, Kossen et al. introduce Semantic Entropy Probes (SEPs) that approximate semantic entropy from a single forward pass (vs. multi-sample SE Farquhar et al. (2024)), achieving strong detection and improving out-of-distribution generalisation Kossen et al. (2024). Clotho also uses the input-side activations, but measures task-specific test adequacy via reference set construction.
@@
Out-of-distribution Detection: For deep neural networks, proposed hidden-state OOD detectors assess how atypical an input is relative to training data. Distance-based methods model class-conditional Gaussians and flag inputs with large Mahalanobis distance Lee et al. (2018). Density-based methods fit mixture models over utterance embeddings, such as GMMs, for unknown intent detection Fan et al. (2020). Reconstruction-based methods rely on hierarchical VAEs Havtorn et al. (2021). Unlike these approaches, which centre on detecting outliers or adversarial cases with respect to training distributions, Clotho addresses task-specific test adequacy by adaptively building a labelled reference set and modelling pre-generation hidden states in LLMs.
9 Conclusion
@@
The data for the studied tasks, as well as the code for replication of the findings of this paper, are publicly available from https://anonymous.4open.science/r/clotho-artifact-DE31.

### v3

@@ 5.3. RQ3: Complementarity Analysis
RQ3 compares Clotho to post-generation approaches. Table 6 begins by comparing the results of correlation analysis, i.e., how well the adequacy scores correlate with pass rates, across Clotho and post-generation metrics: the highest are typeset in bold, and the second highest underlined. Post-generation variance-based metrics generally perform well, with LOHSvar achieving the strongest correlations overall (6/8 for Llama, 5/8 for Mistral). In contrast, token-based confidence measures generally show only moderate correlations across all configurations. Overall, Clotho demonstrates strong performance, despite lacking access to the actual model outputs. Notably, Clotho achieves the best performance on specific tasks, particularly for SYN-BUG, where post-generation approaches typically underperform. We also note that the performance distribution varies significantly across model architectures. For Gemma, Clotho achieves comparable performance to variance-based measures, suggesting particular compatibility with certain model families.
Importantly, these results highlight that Clotho can achieve competitive performance without requiring full output generation, leading to significant efficiency gains. Extracting LIHS incurs negligible additional cost compared to standard inference, taking only a few minutes versus several days for full generation across a test suite. The training cost of the GMM is also minimal, with fitting on 500 samples taking less than 14 seconds (4.4 seconds on average), including time for hyperparameter adaptation.
@@
Figure 7. 3-way Venn diagrams drawn from samples with high adequacy scores (top 25%) each from Clotho, LOHSvar, and TOKent among high-failure-rate inputs ($p_{f}\geq 0.8$) in Llama.
@@
Table 7. Spearman’s rank correlation between scores obtained from OLMs with reference set size 500 and actual pass rates of proprietary LLMs.
Gemma 2 9B Llama 3.1 8B Mistral 7B
@@
5.4. RQ4: Transferability to Proprietary LLMs
Figure 8. Average percentage of detected failures when scores from Open-weight Language Models (OLMs) are applied to proprietary LLMs.
The final research question focuses on the more practical applicability of Clotho: by design, Clotho requires internal representations of the target LLMs that are only available from Open-weight Language Models (OLMs). However, some application may have to rely on stronger, proprietary closed-weight models. Table 7 reports score correlation analysis results, in which the adequacy scores are computed by Clotho using OLMs, while the pass rates are obtained from 2,000 queries made to each of Claude (haiku), GPT (4o-mini), and Gemini (2.5-flash-lite). The results show that the adequacy scores predicted using OLMs can still produce moderate correlations with the actual pass rates of proprietary models: in some cases, they even outperform correlations with pass rates from the OLM itself (contained in the SELF column). For example, Gemma scores achieve higher correlation when applied to Gemini for MODEL-EX than when applied to itself; similarly, Mistral scores perform better when applied to GPT than itself for POS-TAG task. This suggests that Clotho can capture certain generalisable aspects of input difficulty that transcend specific model architectures.
We subsequently evaluate the effectiveness of these transferred adequacy scores for prioritising failure-inducing inputs for the proprietary models. Figure 8 shows the cumulative failure rates for the transferred input orderings for GH-TYPO, JSON-FIX, SPELL-CHK, SYN-BUG, and ODD-ADD for a subset of transfer pairs (best transfer pair for each task): the solid lines denote the average from 10 runs of Clotho, and the coloured bands denote the standard deviation. We also report the Average Percentage of Fault Detected (APFD) (Elbaum et al., 2000), which is essentially the area under curve. Given that average APFD would converge to 0.5 as the number of runs increases, the transferred ordering can achieve significantly effective test prioritisation. This is because proprietary models tend to have lower failure rate. As such, prioritising failure-inducing inputs early on can have bigger impact in proprietary models than in OLMs. Table 8 contains the failure@$N$ from transferred Clotho scores at NN=500, as well as random orderings, per task: the large gains over random ordering aligns with high APFDs we observe in Figure 8. Aggregated across tasks and models, Clotho finds on average 23.75 additional failures at NN=100 and 61.15 at NN=500 over random selection, corresponding to 126.8% and 64.5% improvements, respectively.
Table 8. Number of detected failing inputs at NN labelling budgets (failure@NN) when scores from OLMs are applied to proprietary model (NN=500).
Claude (haiku) GPT (4o-mini) Gemini (2.5-flash-lite)
Gemma 2 9B Llama 3.1 8B Mistral 7B Gemma 2 9B Llama 3.1 8B Mistral 7B Gemma 2 9B Llama 3.1 8B Mistral 7B
@@ 6. Discussion
6.1. Distinguishing Different Failure Modes
Our evaluation focuses on Clotho’s ability to identify inputs that trigger failures, aligning with the core aim of robustness testing: prioritising failure-inducing cases. Since failures are relatively common in LLMs (as confirmed in our studied tasks), testers may also want to distinguish among types of failures. For example, in syntactic bug detection, it is often more valuable to gather one instance from several bug categories (e.g., missing brackets, incorrect keywords) than to collect many instances of the same type.
Since Clotho partitions the input space with a GMM, its components may reveal distinct modes of failure. Targeting low-likelihood inputs from different components could increase the diversity of failures found. Beyond this, building models of failing inputs, potentially one per failure category, alongside the passing-input model offers another path to sharpen failure discovery and characterisation.
@@ 6.1. Distinguishing Different Failure Modes
6.2. Beyond Binary Correctness and Simple Prompts
Tasks in our study, like many existing LLM benchmarks, adopt binary correctness criteria. Yet real-world applications (e.g., counselling chatbot, AI language tutor) often demand richer evaluation, with graded or multi-dimensional measures, such as tone appropriateness, factual accuracy, or consistency. Recent evaluation frameworks (PromptFoo, 2024; AI, 2024a, b; LangChain, 2024) support such score-based assessments, typically assigning continuous scores (e.g., 0-1) using LLM-as-a-judge or human annotators.
We account for non-determinism through majority-pass criterion over repeated runs, and this idea of setting a passing boundary for reference inputs can naturally extend to thresholding the continuous scores. Here, two design choices emerge: (i) aggregating scores across multiple dimensions to construct a unified reference set that satisfy overall quality requirements, or (ii) constructing separate reference sets for each evaluation criterion (e.g., tone-appropriate vs., factually-accurate passing sets). This perspective raises an interesting research question: whether transformer internal states encode signals that distinguish between a certain property-following and property-violating behaviours. If such signals exist, input adequacy could be redefined to align with finer-grained testing goals beyond binary correctness.
@@ 6.2. Beyond Binary Correctness and Simple Prompts
This paper studies the input adequacy for LLMs in the context of relatively simple tasks and prompts, as we believe such prompts are building blocks of more complicated LLM based systems. Future work can consider prompts designed for longer inferences and trajectories, including various agents designed for software engineering tasks (Kang et al., 2024; Zhang et al., 2024b; Bouzenia et al., 2025).
6.3. Alternative Layer Depths
Within the scope of our study, we selected a middle-to-late layer (2/3 depth) guided by the separability observed in the t-SNE visualisation (Section 2.2). To further validate this choice, we additionally evaluate representations from earlier layers (1/3 and 1/2 depths) under the same labelling budget. Overall, we observe that performance degrades as we move away from the selected depth, with the correlation decreasing by 38.8% at 1/3 depth and by 12.5% at 1/2 depth on average. Looking more closely at individual models, we also find that the effective layer depth can vary: while Gemma tends to perform better at deeper layers, Llama and Mistral exhibit comparable performance even at half depth.
These results provide partial support for our initial assumption that middle-to-late layers yield more informative signals for difficulty estimation. At the same time, they suggest that further analysis of layer selection across different models and tasks may be beneficial, potentially requiring additional tuning or model-specific investigation.
7. Threats to Validity
@@
Internal validity could be threatened because LLMs are not fully interpretable, meaning that Clotho might capture latent patterns that correlate with but are not identical to task-specific difficulty; using open-weight models may improve transparency and reproducibility but does not resolve this gap. External validity could also be at risk since Clotho learns task- and model-specific input distributions, which might limit generalisation, especially if other architectures such as reasoning LLMs (e.g., DeepSeek R1 (Guo et al., 2025)) produce different latent spaces; we attempt to reduce this risk by evaluating multiple models and tasks, though broader studies will be useful. Construct validity could likewise be threatened if our chosen metrics do not fully reflect adequacy, and while we use established measures (ROC-AUC, APFD) and a scrutinised, open-source GMM implementation (Pedregosa et al., 2011), these proxies may not perfectly represent the underlying concept.
8. Related Work
@@
Post-Generation Uncertainty: LLM outputs have been widely studied to measure uncertainty and, consequently, detect hallucinations. Chen et al. propose EigenScore, which leverages the eigenvalue decomposition of activation covariance matrices to detect hallucinations  (Chen et al., 2024). The MIND approach of Su et al. goes further and toward online use, showing that unsupervised detectors over streaming internal states can flag hallucinations in real time during decoding (Su et al., 2024). Factoscope from He et al. aggregates multi-layer inner states with a Siamese network architecture to separate factual from non-factual outputs across models, reinforcing that hidden-state geometry correlates with veracity (He et al., 2024). Clotho, on the other hand, uses only the internal hidden space of inputs.
@@
Input-side Internal Activation: Recent work show that an LLM’s input-side internal activations already contain strong signals about errors in the output. Azaria and Mitchell train a lightweight classifier on hidden states to distinguish true statements from the false, outperforming probability-only baselines and demonstrating that pre-generation representations encode truthfulness cues (Azaria and Mitchell, 2023). Snyder et al. probe tokens and activations before the hallucination begins and train binary detectors that can forecast hallucinations in factual QA (Snyder et al., 2024). Extending this, Kossen et al. (Kossen et al., 2024) introduce Semantic Entropy Probes (SEPs) that approximate semantic entropy from a single forward pass (vs. multi-sample SE (Farquhar et al., 2024)), achieving strong detection and improving out-of-distribution generalisation. Clotho also uses the input-side activations, but measures task-specific test adequacy via reference set construction.
@@
Out-of-distribution Detection: For deep neural networks, proposed hidden-state OOD detectors assess how atypical an input is relative to training data. Distance-based methods model class-conditional Gaussians and flag inputs with large Mahalanobis distance (Lee et al., 2018). Density-based methods fit mixture models over utterance embeddings, such as GMMs, for unknown intent detection (Fan et al., 2020). Reconstruction-based methods rely on hierarchical VAEs (Havtorn et al., 2021). Unlike these approaches, which centre on detecting outliers or adversarial cases with respect to training distributions, Clotho addresses task-specific test adequacy by adaptively building a labelled reference set and modelling pre-generation hidden states in LLMs.
9. Conclusion
@@
The data for the studied tasks, as well as the code for replication of the findings of this paper, are publicly available from GitHub [^7] https://github.com/coinse/clotho and Zenodo (Yoon et al., 2026).
Acknowledgments
Juyeon Yoon, Somin Kim, and Shin Yoo are supported by the National Research Foundation of Korea grant (RS-2026-25470171), the Engineering Research Center Program (RS-2021-NR060080), and the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant (RS-2022-II220995), all funded by the Korean Government (MSIT). Robert Feldt is supported by Wallenberg AI, Autonomous Systems, and Software Program (WASP, BoundMiner project), the Swedish Science Foundation (No. 2020-05272, AQUAS) and the Chalmer’s Foundation’s Academic Excellence Program (CAEP-2026).